OpenStudy (anonymous):

Related Rates Calculus Question A square plate is cooling and contracting at a rate of 3cm^2/sec. How fast is the length of an edge decreasing when the area of the plate is 12cm^2?

OpenStudy (abb0t):

do they give you an equation to use? I think area of a square is A = xy? height x width or length x width? I forget. Use that equation and differentiate the whole thing w/ respect to time. If it is A = xy differentiate to get dA/dt = x(dx/dt) + y(dy/dt) <--- product rule plug in your knowns, use algebra to rearrange and find your unknown.

OpenStudy (anonymous):

I believe that the area of a square is A=S^2? because squares have the same sides throughout? im not really sure

OpenStudy (abb0t):

Differentiate to get: \[\frac{ dA }{ dt }= 2s(\frac{ ds }{ dt })\]

OpenStudy (abb0t):

Sorry im getting myself confused now. When I see cooling, I think of newtons law of cooling, partial differentials.

OpenStudy (anonymous):

oh, its ok, i am solving it right now, may you please check my work afterwards? thank you :)

OpenStudy (abb0t):

fa sho

OpenStudy (anonymous):

oh wait if the area of the plane is 12cm^2, how do I find the sides?

OpenStudy (abb0t):

2 sides means you multiply it by two. Assuming they are of equal length.

OpenStudy (anonymous):

so \[\sqrt{12}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Area of Square: A = x² = √ 12 -> x = 2√3 Thus: -3 = 2 * 2√3 dx/dt Can you solve for dx/dt now?

OpenStudy (anonymous):

hahaha! i was wrong its -sqrt3/4

OpenStudy (anonymous):

thank you :)